Stability of the Absolutely Continuous Spectrum of Random Schrödinger Operators on Tree Graphs

Robert Sims University of California at Davis, USA

The objective of this work is to present results on the stability of the absolutely continuous spectrum of Schrodinger operators on tree graphs, under the addition of weak but extensive disorder in the form of a random potential.